OpenStudy (anonymous):

(2/3)x - (5/6)x + 7 = 1/15(33-5x) - 6/5

OpenStudy (lucaz):

first calculate (1/15)*(33-5x)

OpenStudy (anonymous):

(( 2 3 )(x)−( 5 6 )(x))+7=( 1 15 )(33−5x)− 6 5 I will try to solve your equation. −0.166667x+7=−0.333333x+1(Simplify both sides of the equation) −0.166667x+7+0.3333333333333333x=−0.333333x+1+0.3333333333333333x(Add 0.3333333333333333x to both sides) 0.166667x+7=1 0.166667x+7+−7=1+−7(Add -7 to both sides) 0.166667x=−6 (0.166667x) 0.16666666666666669 = (−6) 0.16666666666666669 (Divide both sides by 0.16666666666666669)

OpenStudy (anonymous):

I won't give the answer. This is how you break it down.

OpenStudy (anonymous):

I got -35.999928

OpenStudy (lucaz):

it would be better if you work with fractions

OpenStudy (anonymous):

thank you

OpenStudy (lucaz):

unless you're allowed to use calculator

OpenStudy (anonymous):

So the answer is : x=−36